OpenStudy (anonymous):

What is the equation of the line that passes through the point 9-1,4) and is perpendicular to the equation of the graph of y=4x-3?

OpenStudy (p0sitr0n):

perpendicular of y=mx+b means its slope is -1/m

OpenStudy (anonymous):

Sorry the 9 is supposed to be an open parenthesis

OpenStudy (p0sitr0n):

so y=-1/4x+b 4=-1/4*-1+b b=3/4 y=-1/4 x + 3/4

OpenStudy (ybarrap):

To find the slope of the line perpendicular to y=4x-3 you need to know the slope of this equation. Let's call it m, so m = 4 The negative reciprocal of this is the slope of the line perpendicular to this equation: m' = -1/4 So our equation that is perpendicular to your equation is y' = m'x + b' Now plug in the point you have to find b' 4=m' (-1) + b' 4=(-1/4)(-1) + b' b' = 4-1/4 So the equation perpendicular to y=4x -3 is y' = (-1/4) x + b' Where b' is given above. Make sense?

OpenStudy (ybarrap):

correction b'=1/4 - 4